# UniGLM: Training One Unified Language Model for Text-Attributed Graphs

### v1

@@ Abstract
Representation learning on text-attributed graphs (TAGs), where nodes are represented by textual descriptions, is crucial for textual and relational knowledge systems and recommendation systems. Currently, state-of-the-art embedding methods for TAGs primarily focus on fine-tuning language models (e.g., BERT) using structure-aware training signals. While effective, these methods are tailored for individual TAG and cannot generalize across various graph scenarios. Given the shared textual space, leveraging multiple TAGs for joint fine-tuning, aligning text and graph structure from different aspects, would be more beneficial. Motivated by this, we introduce a novel Unified Graph Language Model (UniGLM) framework, the first graph embedding model that generalizes well to both in-domain and cross-domain TAGs. Specifically, UniGLM is trained over multiple TAGs with different domains and scales using self-supervised contrastive learning. UniGLM includes an adaptive positive sample selection technique for identifying structurally similar nodes and a lazy contrastive module that is devised to accelerate training by minimizing repetitive encoding calculations. Extensive empirical results across 9 benchmark TAGs demonstrate UniGLM’s efficacy against leading embedding baselines in terms of generalization (various downstream tasks and backbones) and transfer learning (in and out of domain scenarios). The code is available at https://github.com/NYUSHCS/UniGLM.
UniGLM: Training One Unified Language Model for
Text-Attributed Graphs
Yi Fang1, Dongzhe Fan1, Sirui Ding2, Ninghao Liu3, Qiaoyu Tan1 1 New York University Shanghai 2 University of California, San Francisco; 3 University of Georgia {yf2722, qiaoyu.tan}@nyu.edu
1 Introduction
@@
Text-attributed graphs (TAGs) have been widely adopted to represent complex relationships between textual entities in real-world textual and relational knowledge systems, including social media, recommendation systems, and knowledge base. Unlike standard graphs, nodes in TAGs are represented by text attributes. A typical example is academic citation network, where nodes represent scientific papers and edges indicate citations. To learn from TAGs, graph embedding (GE) Perozzi et al. (2014); Grover and Leskovec (2016); Zhang et al. (2018); Wu et al. (2020), which maps nodes into embedding vectors that preserve both textual and structure information, has recently garnered significant attention.
@@
Prior GE studies He et al. (2024); Chen et al. (2024) on TAGs primarily focus on two stages: text transformation and graph structure modeling. In the first stage, text attributes are transformed into numerical feature vectors via shallow embedding models such as Word2vec Mikolov et al. (2013) and Bag-of-Words (BoW) Harris (1954). Subsequently, the transformed node features, along with graph structure, are often fed into graph neural networks (GNNs) Kipf and Welling (2016a); Zhou et al. (2020) for structural analysis. Although these methods are straightforward, they may be suboptimal in effectively integrating text semantics and structure knowledge.
@@
In recent years, there has been a notable shift in interest from shallow models to pretrained language models (PLMs) such as BERT Devlin et al. (2019). The high-level idea is to jointly learn text knowledge and graph structure within a single encoder, either by developing nested graph-BERT architectures Yang et al. (2021) or by designing structure-aware training signals Chien et al. (2022); Jin et al. (2023). Despite their popularity, these methods face limitations in generalization capability because they fine-tune the BERT model for a single particular TAG, making it ineffective for transferring to other TAGs for representation learning. Given that text attributes provide a unified semantic space across different TAGs, leveraging multiple TAGs for a joint fine-tuning is a promising yet under-explored research direction, supported by the scaling law Kaplan et al. (2020).
@@
However, training a unified BERT model for multiple TAGs presents several challenges. Firstly, extracting effective structural information across various graph scenarios while maintaining their unique statistics for LM fine-tuning is difficult. Given the diversity and variability of TAGs, local structures such as node degrees and global structures within the graph vary from nodes to nodes and graphs to graphs. Secondly, directly combining multiple TAGs for joint BERT training may suffer from memory and training efficiency issues due to the non i.i.d. nature of graphs. Unlike pure text-based LM training, textual nodes in TAGs are strongly correlated with one another. Consequently, anchor nodes and their structurally similarly neighbors need to be processed by BERT simultaneously, leading to significant trade-offs in computational and memory consumption.
@@
To address the aforementioned challenges, we propose a novel unified graph language model (UniGLM) framework, the first self-supervised language model pre-training method tailored for multiple TAGs. The key idea is to enhance language model’s (e.g., BERT) graph embedding capability by fine-tuning it using large scale, diverse and cross-domain text-to-structure knowledge based on contrastive learning. Specifically, to tackle the first challenge, we introduce an adaptive positive sample selection technique that identifies positive samples by considering each node’s local, global, and graph-specific contexts. This sampling strategy is personalized and can effectively align textual nodes and their important neighbors well across different TAGs. To address the second challenge, we devise a dynamic memory bank to encode positive samples off-the-fly, thereby accelerating the training speed by avoiding repetitive encoding of positive samples’ text attributes via BERT. Our major contributions are summarized below.
@@
  • We propose an adaptive positive sample selection method for sampling positive samples of each node for contrastive learning. Unlike standard sampling strategies, our personalized scheme identifies positive samples based on nodes’ local, global, and graph-related contexts, thereby unifying graph structures across various TAGs.
  • We devise a simple yet effective dynamic embedding table scheme to encode sampled positive samples off-the-fly during mini-batch training. By maintaining an external memory bank to update and retrieval embeddings of positives examples, we accelerate the training process using historical embeddings as supervision.
@@
Figure 1: The proposed UniGLM framework. The UniGLM framework trains a unified graph encoder across multiple TAGs using contrastive learning, instead of learning separate language models for each TAG. To ensure effective and efficient textual-to-structure alignment, we introduce an adaptive positive sample selection scheme and a lazy contrastive strategy. UniGLM serves as a foundational embedding model for TAGs, consistently delivering strong performance across various downstream tasks and backbones.
2 Preliminaries
@@
In this section, we introduce notations, formulate the research problem, and illustrate the motivation behind learning from multiple TAGs.
@@
Notations and Problem Formulation. We are given $m$ TAGs, denoted as $\{G_{i}|i=1,2,...,m\}$, where $\mathcal{G}_{i}=(\mathcal{V},\mathcal{T},\mathbf{A})$ represents the $i$-th TAG with $n_{i}$ nodes. $\mathcal{V}$ is the set of nodes and $\mathbf{A}\in\mathbb{R}^{n_{i}\times n_{i}}$ is the adjacency matrix. Each node $v\in\mathcal{V}$ is associated with a textual attribute $\mathcal{T}_{v}$, and $\mathcal{T}=\{\mathcal{T}_{v}\mid v\in\mathcal{V}\}$ represents the set of these attributes.
@@
TAG Embedding. Given a TAG $G_{i}=(\mathcal{V},\mathcal{T},\mathbf{A})$, a standard embedding model aims to learn a graph encoder $f_{i}$ that maps nodes in $G_{i}$ into embedding vectors, preserving both textual ($\mathcal{T}$) and structure ($\mathbf{A}$) knowledge. Therefore, for $m$ TAGs, traditional methods will learn $m$ independent graph encoders, denoted as $\{f_{i}\}_{i=1}^{m}$.
@@
Motivation. Learning one graph encoder for each particular TAG is the de facto standard in state-of-the-art graph embedding literature. However, we argue that this setting is suboptimal for two main reasons. i) Deployment inefficiency. As discussed above, this learning procedure requires the development of $m$ separate graph encoders for all TAGs, significantly increasing the deployment and maintenance costs in practice. ii) Limited performance. Given the shared textual space across various TAGs, pre-training a language model for a single TAG is inherently less effective because it cannot leverage the text-to-structural knowledge across TAGs. According to the scaling law Kaplan et al. (2020), incorporating more structure-aware textual knowledge for collaborative language model fine-tuning may be advantageous.
Motivated by this, we aim to explore learning from multiple TAGs, as follows.
@@
3 The Proposed Method
In this section, we present the details of UniGLM, as depicted in Figure 1. First, we introduce the contrastive-based collective language model pre-training pipeline in Section 3.1. Then, in Section 3.2, we elaborate on the methodology to adaptively select positive samplesfor collaborative training. Finally, in Section 3.3, we introduce a simple but effective optimization strategy (Embedding Table) to accelerate our learning process.
3.1 The Overall Pipeline of Collaborative Language Model Pre-training
Learning from multiple TAGs is challenging due to the heterogeneity of textual attributes and graph structures. Existing methods address this challenge by either employing GNN-nested transformer, as seen in Graphformers Yang et al. (2021) to capture both textual attributes and their correlations among nodes, or by adopting structure-aware objectives to fine-tune LMs Chien et al. (2022). While the former approach is effective, it may encounter efficiency issues due to the combination of GNNs and Transformers. Conversely, the latter approach is computationally efficient, but it predominantly focuses on learning from individual TAGs, leaving joint learning from diverse TAGs relatively under-explored.
@@ 3.1 The Overall Pipeline of Collaborative Language Model Pre-training
$\begin{split}\mathcal{L}_{v}=-\sum_{u\in\mathcal{S}_{v}}\log\frac{\exp\left({\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{u})})/{\tau}\right)}{\sum\limits_{k\in\mathcal{B}}\exp\left({\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{k})})/{\tau}\right)},\end{split}$ (1)
@@
where $\mathcal{S}_{v}$ denotes the set of structurally similar nodes to $v$, $\mathcal{B}$ is the sampled batch set in mini-batch training with $v\in\mathcal{B}$, and $\tau$ signifies the temperature parameter. $\text{sim}(\cdot,\cdot)$ is a similarity function such as the inner product, and $f$ is implemented as BERT by default, enabling the encoding of text attributes into embedding vectors. By optimizing Eq. (1), the language encoder $f$ is trained to generate similar representations for node $v$ and nodes in $\mathcal{S}_{v}$, while simultaneously pushing away the representations of $v$ and other nodes in the mini-batch set $\mathcal{B}$. Notably, as nodes in $\mathcal{B}$ are randomly sampled across $m$ TAGs, Eq. (1) provides a simple yet effective way to learn from nodes in various scenarios due to its instance-wise discriminator nature.
@@
While conceptually simple and feasible, learning through Eq. (1) faces two major challenges in practice. C1: The structurally similar node set $\mathcal{S}_{v}$ of $v$ is not well-defined. Given the heterogeneity of various TAGs and node-level statistics, random sampling based on neighbors may be suboptimal for capturing the diversity between nodes across different TAGs. C2: It presents a trade-off between model performance and training efficiency. The computational costs of Eq. (1) are determined by the number of positive samples ($|\mathcal{S}_{v}|$) and batch size ($|\mathcal{B}|$), as it requires the language model to encode $(|\mathcal{S}_{v}|+1)|\mathcal{B}|$ text sequences per iteration. While reducing the number of positive samples can accelerate training, it may degrade performance since sufficient structural information is critical for graph contrastive learning You et al. (2020); Zhu et al. (2021); Zhang et al. (2024). In Section 3.2 and Section 3.3, we introduce two strategies to address these challenges, respectively.
3.2 Adaptive Positive Sample Selection

### v2

# UniGLM: Training One Unified Language Model for Text-Attributed Graph Embedding

@@ Abstract.
Representation learning on text-attributed graphs (TAGs), where nodes are associated with textual descriptions, is crucial for textual and relational knowledge systems, such as social media and recommendation scenarios. However, state-of-the-art embedding methods for TAGs primarily focus on fine-tuning pre-trained language models (PLMs) using structure-aware training objectives. While effective, these methods are tailored for individual TAG and cannot generalize across various graph scenarios. Given the shared textual space, leveraging multiple TAGs for joint fine-tuning, aligning text and graph structure from different aspects, would be more beneficial. Therefore, we propose the Unified Graph Language Model (UniGLM), a novel foundation model pretrained over multiple TAGs from a variety of domains, which can generalize well to both in-domain and cross-domain graph scenarios. Specifically, UniGLM fine-tunes well-established PLMs (e.g., Sentence-BERT) using a domain-aware contrastive learning objective that unifies structure heterogeneity and node statistics across various domains with an adaptive and learnable positive sample selection scheme. Additionally, a lazy updating module is introduced to speed up training by reducing repetitive encoding of positive samples. Extensive datasets across multiple domains, downstream tasks (node classification and link prediction), and a spectrum of graph backbones (supervised and self-supervised graph models) are conducted to compare UniGLM with state-of-the-art baselines. Our empirical observations suggest that UniGLM can generate informative representations for cross-domain graphs observed in the training. More importantly, UniGLM also exhibits competitive transfer ability in encoding unseen TAGs that are not used for training. This study provides deep insights into how to adapt PLMs to graph data and demonstrates the potential of building foundation model for graph representation learning.
Graph embedding, Language model, Text-attributed graph
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germany.booktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanyisbn: 979-8-4007-1329-3/25/03doi: 10.1145/3701551.3703586
1. Introduction
@@
Text-attributed graphs (TAGs) have been widely adopted to represent complex relationships between textual entities in real-world textual and relational knowledge systems, including social media, recommendation systems, and knowledge base. Unlike standard graphs, nodes in TAGs are represented by text attributes. A typical example is academic citation network, where nodes represent scientific papers and edges indicate citations. To learn from TAGs, graph embedding (GE) (Perozzi et al., 2014; Grover and Leskovec, 2016; Zhang et al., 2018; Wu et al., 2020), which maps nodes into embedding vectors that preserve both textual and structure information, has recently garnered significant attention.
@@
Prior GE studies (He et al., 2024; Chen et al., 2024a) on TAGs primarily focus on two stages: text transformation and graph structure modeling. In the first stage, text attributes are transformed into numerical feature vectors via shallow embedding models such as Word2vec (Mikolov et al., 2013) and Bag-of-Words (BoW) (Harris, 1954). Subsequently, the transformed node features, along with graph structure, are often fed into graph neural networks (GNNs) (Kipf and Welling, 2016a; Zhou et al., 2020) for representation learning, where the objective is to either (1) reconstruct the observed links (Kipf and Welling, 2016b; Li et al., 2023; Tan et al., 2023) or transformed features (Hou et al., 2022; Shi et al., 2023; Hou et al., 2023), or (2) learn invariant representations between graph augmentations (You et al., 2020; Zhang et al., 2024; Zhu et al., 2021; Xu et al., 2021; Thakoor et al., [n. d.]), i.e., graph contrastive learning. Despite their promising results, they may be suboptimal in effectively integrating text semantics and structure knowledge.
@@
In recent years, there has been a notable shift of interest in graph embedding from shallow models to pre-trained language models (PLMs) such as BERT (Devlin et al., 2019) and Sentence-Bert (Reimers and Gurevych, 2019). The high-level idea is to jointly learn text knowledge and graph structure within a single encoder, either by developing nested graph-BERT architectures (Yang et al., 2021) or by designing structure-aware training signals (Chien et al., 2022; Jin et al., 2023). Despite their popularity, these methods face limitations in generalization capability because they fine-tune the PLM model for a single particular TAG, making it ineffective for transferring to other TAGs for graph embedding. Given that text attributes provide a unified semantic space across different TAGs, leveraging multiple TAGs for a joint fine-tuning is a promising yet under-explored research direction, supported by the scaling law (Kaplan et al., 2020).
@@
However, training a unified PLM model for multiple TAGs presents several challenges. Firstly, extracting effective structural information across various graph scenarios while maintaining their unique statistics for LM fine-tuning is difficult. Given the diversity and variability of TAGs, local structures such as node degrees and global structures within the graph vary from nodes to nodes and graphs to graphs. Secondly, directly combining multiple TAGs for joint PLM training may suffer from memory and training efficiency issues due to the non-i.i.d. nature of graphs. Unlike pure text-based LM training, textual nodes in TAGs are strongly correlated with each other. Consequently, anchor nodes and their structurally similar neighbors need to be processed by PLM simultaneously, leading to significant trade-offs in computational and memory consumption.
@@
To address the aforementioned challenges, we propose a novel unified graph language model (UniGLM) framework, a pure PLM-based graph embedding approach tailored for multiple TAGs. The key insight is to enhance PLM’s (e.g., Sentence-BERT) graph embedding capability by fine-tuning it using large-scale, diverse, and cross-domain text-to-structure knowledge based on domain-aware contrastive learning. Specifically, to tackle the first challenge, we introduce an adaptive and learnable positive sample selection technique that identifies positive samples of an anchor node considering its local, global, and graph-specific contexts. This sampling strategy is personalized and can effectively align textual nodes and their important neighbors well across different TAGs. After that, the PLM model is optimized by maximizing mutual information between the anchor node representation and the most informative positive sample adaptively selected from the candidate pool in a learnable fashion. This learnable positive generation scheme further enhances the learning capacity of UniGLM by focusing on positive samples that provide sufficient permutations to contrast anchor nodes. To address the second challenge, we devise a dynamic memory bank to encode positive samples off-the-fly, thereby accelerating the training speed by avoiding repetitive encoding of positive samples’ text attributes via PLM. The key contributions of our work are summarized as follows:
@@
  • We propose an adaptive and learnable positive sample selection method for sampling positive samples of each node for domain-aware contrastive learning. Unlike standard sampling strategies, our personalized scheme identifies positive samples based on nodes’ local, global, and graph-related contexts, thereby unifying graph structures across various TAGs.
  • Subsequently, a learnable positive generation module is introduced to adaptively select positive samples from the pool aimed at enhancing the contrastive objective by focusing on informative positives (a.k.a. effective node augmentation). Besides, we devise a simple yet effective dynamic embedding table scheme to encode sampled positive samples off-the-fly, accelerating the training process by using historical knowledge as supervision.
@@
2. Related Work
2.1. Representation Learning on TAGs
Text-attributed graphs (TAGs) have become a pivotal area of study due to their ability to encapsulate complex relationships between textual data and graph structures. Initially, TAGs employed shallow embedding techniques that were limited in their capacity to meld textual content robustly with graph topology. With the advent of pre-trained language models (PLMs), a more sophisticated extraction of text features became possible. These features are then integrated into graph neural networks (GNNs). However, challenges persist as the integration often leads to suboptimal performance. To address these issues, recent methodologies have sought to more effectively amalgamate PLM features with GNN architectures, thereby enhancing overall model effectiveness. Notably, LM-GNN(Ioannidis et al., 2022) has pioneered the joint training of large-scale language models with GNNs, creating a unified learning framework. GIANT(Chien et al., 2022) leverages extreme multi-label classification to fine-tune language models using graph-based data. Furthermore, GLEM(Zhao et al., 2023b) introduces an innovative iterative update mechanism between LM and GNN components in its Expectation-Maximization framework. Another significant advancement is Patton(Jin et al., 2023), which introduces a graph-aware pre-training approach that includes network-contextualized masked language modeling and masked node prediction, specifically designed to enhance the relevance of pre-trained embeddings for graph-oriented tasks.
2.2. Graph Embedding Foundation Models
Graph Foundation Models (GFMs) are designed to operate across a diverse range of graphs and tasks, marking a significant step toward universal graph embedding. OpenGraph(Xia et al., 2023), proposed by Xia et al., excels in zero-shot learning scenarios, illustrating the model’s ability to generalize well without tailored training. Self-supervised learning (SSL) techniques play a critical role in the pre-training of GFMs, with Zhao et al.(Zhao et al., 2023a) classifying SSL tasks to better harness graph-embedded knowledge, thus enhancing model adaptability. Liu et al.(Liu et al., 2023) discuss the ongoing challenges and potential future directions for GFMs, emphasizing the need for continued innovation. Meanwhile, Tan et al.(Tan et al., 2023) introduce an advanced structure reconstruction approach in their model, aiming to further boost the generalizability of GFMs. For those interested in a deeper dive into Graph Foundation Models, particularly regarding their generalization capabilities, scalability, and efficiency, the comprehensive analysis provided in the paper by Xu et al.(Xu et al., 2024) is recommended. This work not only discusses various self-supervised GNN models but also includes experimental code, facilitating hands-on exploration and further research in this rapidly evolving field.
2.3. Graph-aware Large Language Models
Beyond the typical encoder-only configurations, decoder-only large language models (LLMs) are now being adapted for graph-related tasks. GraphGPT (Tang et al., 2024) employs a novel projector mechanism to align graph tokens with a pre-trained vocabulary through instruction tuning, facilitating better model understanding and generation of graph structures. LLaGA (Chen et al., 2024b) melds the capabilities of large language models (LLMs) with the unique requirements of graph-structured data, preserving the versatility of LLMs while reformulating graph data to be more amenable to standard LLM inputs. Additionally, MolecularGPT(Liu et al., 2024) showcases the potential of LLMs in acting as efficient few-shot predictors for molecular properties, highlighting the model’s adaptability to specialized domains.
3. Preliminary
@@
Figure 1. The proposed UniGLM framework. The UniGLM framework trains a unified graph encoder across multiple TAGs using domain-aware contrastive learning, instead of learning separate language models for each TAG. To ensure effective and efficient textual-to-structure alignment, we introduce an adaptive and learnable positive sample selection scheme and a lazy updating strategy. UniGLM serves as a foundational embedding model for TAGs, consistently delivering strong performance across various downstream tasks and backbones.
@@
In this section, we introduce the notation, research problem definition, and the motivation behind learning from multiple TAGs.
@@
Notations and Problem Definition. We are given $m$ TAGs, denoted as $\{G_{i}|i=1,2,...,m\}$, where $G_{i}=(\mathcal{V},\mathcal{T},\mathbf{A})$ represents the $i$-th TAG with $n_{i}$ nodes. $\mathcal{V}$ is the set of nodes and $\mathbf{A}\in\mathbb{R}^{n_{i}\times n_{i}}$ is the adjacency matrix. Each node $v\in\mathcal{V}$ is associated with a textual attribute $\mathcal{T}_{v}$, and $\mathcal{T}=\{\mathcal{T}_{v}\mid v\in\mathcal{V}\}$ denotes the attribute set.
@@
TAG Embedding. Given a TAG $G_{i}=(\mathcal{V},\mathcal{T},\mathbf{A})$, a standard embedding model aims to learn a graph encoder $f_{i}$ that maps nodes in $G_{i}$ into embedding vectors, preserving both textual ($\mathcal{T}$) and structure ($\mathbf{A}$) knowledge. Therefore, for $m$ TAGs, traditional methods will learn independent $m$ graph encoders, denoted as $\{f_{i}\}_{i=1}^{m}$.
@@
Motivation. Learning one graph encoder for each particular TAG is the de facto standard in state-of-the-art graph embedding literature. However, we argue that this setting is suboptimal for two main reasons. i) Deployment inefficiency. As discussed above, this learning procedure requires the development of $m$ separate graph encoders for all TAGs, significantly increasing the deployment and maintenance costs in practice. ii) Limited performance. Given the shared textual space across various TAGs, pre-training a language model for a single TAG is inherently less effective because it cannot leverage the text-to-structural knowledge across TAGs. According to the scaling law (Kaplan et al., 2020), incorporating more structure-aware textual knowledge for collaborative language model fine-tuning may be advantageous. Motivated by this, we aim to explore learning from multiple TAGs, as follows.
@@
4. Methodology
In this section, we present the details of UniGLM, as depicted in Figure 1. First, we introduce the contrastive-based collective language model pre-training pipeline in Section 4.1. Then, in Section 4.2, we elaborate on the methodology to adaptively select positive samples for collaborative training. After that, we introduce how to aggregate positive samples in Section 4.2.2. Finally, in Section 4.3, we introduce a simple but effective optimization strategy (Embedding Table) to accelerate our learning process.
4.1. The Overall Pipeline of Collaborative Language Model Pre-training
Learning from multiple TAGs is challenging due to the heterogeneity of textual attributes and graph structures. Existing methods address this challenge by employing a GNN-nested transformer, as seen in Graphformers (Yang et al., 2021) to capture both textual attributes and their correlations between nodes, or by adopting structure-aware objectives to fine-tune LMs (Chien et al., 2022). While the former approach is effective, it may encounter efficiency issues due to the combination of GNNs and Transformers. Conversely, the latter approach is computationally efficient, but it predominantly focuses on learning from individual TAGs, leaving joint learning from diverse TAGs relatively under-explored.
@@ 4.1. The Overall Pipeline of Collaborative Language Model Pre-training
(1) $\begin{split}\mathcal{L}_{v}=-\sum_{u\in\mathcal{S}_{v}}\log\frac{\exp\left({\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{u})})/{\tau}\right)}{\sum\limits_{k\in\mathcal{B}}\exp\left({\text{sim}(f(\mathcal{T}_{v}),f(\mathcal{T}_{k})})/{\tau}\right)},\end{split}$
@@
where $\mathcal{S}_{v}$ denotes the set of structurally similar nodes to $v$, $\mathcal{B}$ is the sampled batch set in mini-batch training with $v\in\mathcal{B}$, and $\tau$ signifies the temperature parameter. $\text{sim}(\cdot,\cdot)$ is a similarity function such as the inner product, and $f$ is a PLM encoder, enabling the encoding of text attributes into embedding vectors. By optimizing Eq. (1), the language encoder $f$ is trained to generate similar representations for node $v$ and nodes in $\mathcal{S}_{v}$, while simultaneously pushing away the representations of $v$ and other nodes in the mini-batch set $\mathcal{B}$. Notably, as nodes in $\mathcal{B}$ are randomly sampled across $m$ TAGs, Eq. (1) provides a simple yet effective way to learn from nodes in various scenarios due to its instance-wise discriminator nature.
@@
While conceptually simple and feasible, learning through Eq. (1) faces two major challenges in practice. C1: The structurally similar node set $\mathcal{S}_{v}$ of $v$ is not well-defined. Given the heterogeneity of various TAGs and node-level statistics, random sampling based on neighbors may be suboptimal for capturing the diversity between nodes across different TAGs. C2: It presents a trade-off between model performance and training efficiency. The computational costs of Eq. (1) are determined by the number of positive samples ($|\mathcal{S}_{v}|$) and batch size ($|\mathcal{B}|$), as it requires the language model to encode $(|\mathcal{S}_{v}|+1)|\mathcal{B}|$ text sequences per iteration. While reducing the number of positive samples can accelerate training, it may degrade performance since sufficient structural information is critical for graph contrastive learning (You et al., 2020; Zhu et al., 2021; Zhang et al., 2024). In Section 4.2 and Section 4.3, we introduce two strategies to address these challenges, respectively.
4.2. Adaptive and Learnable Positive Scheme